OpenStudy (caozeyuan):

vector question The line l1 passes through the point A whose position vector is 3i+j+2k and is parallel to the vector i+j The line l2 pass the through point B whose position vector is -i-k and is parallel to the vector j+2k . The point P is on l1 and point Q is on l2 and PQ is perpendicular to both l1 and l2 . FInd the position vector of Q

OpenStudy (caozeyuan):

@amistre64

OpenStudy (amistre64):

position vector of Q?

OpenStudy (amistre64):

The line l1 passes through the point A whose position vector is 3i+j+2k and is parallel (has a direction) to the vector i+j L1: x=3+1t y=1+1t z=2+0t --------------------------- The line l2 pass the through point B whose position vector is -i-k and is parallel to the vector j+2k . L2: x=-1+0s y= 0+1s z=-1+2s ----------------------- The point P is on l1 point Q is on l2 PQ is perpendicular to both l1 and l2 . cross product FInd the position vector of Q

OpenStudy (amistre64):

the smallest distance between L1 and L2 should define the points of P and Q

OpenStudy (caozeyuan):

In the previous question I find the shortest distance to be 3

OpenStudy (amistre64):

when you know the shortest distance, then you need to be able to place it as well

OpenStudy (amistre64):